The sum of two digits number is 12. The tens digit in two more than the unit digit . Find the number

Found 3 solutions by ikleyn, macston, fractalier:
Answer by ikleyn(52802)

.
75.

You can quickly try all the numbers from 1 to 9 with the sum of 12.

It is faster than to write and to solve an equation.


Answer by macston(5194)

.
N=units digit; T=tens digit
.
T=N+2
.
T+N=12
(N-2)+N=12
2N=10
N=5
ANSWER 1: The units digit is 5.
.
T=N+2=5+2=7
ANSWER 2: The tens digit is 7
.
ANSWER: The number is 75.
.

Answer by fractalier(6550)

One can guess at it, but if we call the tens digit x and the ones digit y we have
x + y = 12
x = y + 2 or x - y = 2
Now add them together and get
x + y = 12
x - y = 2
------------
2x = 14
x = 7
y = 5
75 is your number.